Consider the American put option with value P(S,t), where S is the asset price and t is time. Let E be the strike price, T be the expiry date, r be the interest rate, σ be the volatility, and Sf(t) be an optimal exercise boundary. Write down the partial differential equation on P(S,t) when S > Sf(t), the formula for P(S,t) when S < Sf(t), the inequality which P(S,t) must satisfy for S < Sf(t), the two conditions at S = Sf(t), the condition at t = T, the boundary condition at S = 0, and the boundary condition at S → ∞


1
Expert's answer
2021-01-19T04:14:10-0500

S > Sf(t):

"∂P +\frac{1}{ 2} σ^2S^2∂^2sP + rS∂sP − rP = 0"


S < Sf(t):

"P (S, T) = max(E \u2212 S, 0)"


for S = Sf (t) and S = ∞:

"P (Sf (t), t) = E \u2212 Sf (t)"


"\u2202SP (Sf (t), t) = \u22121"


"P (+\u221e, t) = 0"